The work function for caesium atom is 1.9eV. calculate the threshold wavelength and the threshold frequency of the radiation. If the caesium element is irradiated with a wavelength 500nm,calculate the kinetic energy and the velocity of the ejected photoelectron.

Grade:11

1 Answers

ahmad areej
5 years ago
threshold frquency = Energy(in Joules)/plank’s constant.
                               =1.9*1.6*10^-19/6.626*10^-34 
                               =4.587*10^14
threshold wavelength = c/threshhold frequency
                                   =3*10^8/4.587*10^14
                                   =6540 * 10^-10
K.E = Energy total – work function
       = (6.626*3*10^-26)/500 *10^-9   – (1.9*1.6*10^-19)
       = 0.335 * 10^-19
       = 335*10^-22
 
 
v^2 = 2*K.E/m
       =2* 335*10^-22/1.67*10^-27
       =670 * 10^5?1.67
       =  67 * 10^6/1.67
v^2 =40.1197 * 10^6
 v   =1290 m/s